# ArchRAG: Attributed Community-based Hierarchical Retrieval-Augmented Generation

### v1

@@ 12      ${G}(V,E)\leftarrow$ build a new graph using $C$ and $E^{\prime}$;
14until $T$=True;
return ${HC}$;
@@
3.1.3. C-HNSW index.
Given a query, to efficiently identify the most relevant information from each layer of the hierarchical tree $\Delta$, a naive method is to build a vector database for the attributed communities in each layer, which is costly in both time and space. To tackle this issue, we propose to build a single hierarchical index for all the communities. Recall that the attributed communities in $\Delta$ form a tree structure, and the number of nodes decreases as the layer level increases. Since this tree structure is similar to the HNSW (Hierarchical Navigable Small World) index which is the most well-known index for efficient ANN search (Malkov and Yashunin, 2018) (Appendix A.2 introduces its details), we propose to map entities and attributed communities of $\Delta$ into high-dimensional nodes, and then build a unified Community-based HNSW (C-HNSW) index for them.
$\bullet$ The structure of C-HNSW. Conceptually, the C-HNSW index is a list of simple graphs with links between them, denoted by $\mathcal{H}=({\mathcal{G}},{L_{inter}})$ with $\mathcal{\mathcal{G}}=\{G_{0}=(V_{0},E_{0}),G_{1}=(V_{1},E_{1}),\cdots,G_{L}=(V_{L},E_{L})\}$, where $G_{i}$ is a simple graph and each node of the simple graph corresponds to an attributed community or entity. The number $L$ of layers (simple graphs) of $\mathcal{H}$ is the same as the that of $\Delta$. Specifically, for each attributed community or entity in the $i$-the layer of $\Delta$, we map it to a high-dimensional node in the $i$-th layer of $\mathcal{H}$ by using a language model (e.g., nomic-embed-text (Nussbaum et al., 2024)).
After obtaining the high-dimensional nodes, we establish two types of links between them, i.e., intra-layer and inter-layer links:
  • Intra-layer links: These links exist between nodes in the same layers. Specifically, for each node in each layer, we link it to at least $M$ nearest neighbors within the same layer, where $M$ is a predefined value, and the nearest neighbors are determined according to a given distance metric $d$. Thus, all the intra-layer links are edges in all the simple graphs:
    (1) $L_{intra}=\bigcup_{i=0}^{L}E_{i}$
  • Inter-layer links: These links cross two adjacent layers. Specifically, we link each node in each layer to its nearest neighbor in the next layer. As a result, all the inter-layer links can be represented as follows:
    (2) $L_{inter}=\bigcup_{i=1}^{L}\{(v,\psi(v))|v\in V_{i},\psi(v)\in V_{i-1}\},$
    where $\psi(\cdot):V_{i}\rightarrow V_{i-1}$ is the injective function that identifies the nearest neighbor of each node in the lower layer.
For example, in Figure 2, the C-HNSW index has three layers (simple graphs), incorporating all the attributed communities. Within each layer, each node is connected to its two nearest neighbors via intra-layer links, denoted by undirected edges. The inter-layer links are represented by arrows, e.g., the green community at layer $L_{1}$ is connected to the green entity (its nearest neighbor at layer $L_{0}$).
@@
$\bullet$ The construction of C-HNSW. A naive approach to build the C-HNSW index is to build nodes first and then establish the two types of links by finding the nearest neighbors of each node. However, the process of finding the nearest neighbor is costly. To accelerate the construction, we propose a top-down approach by borrowing the idea of HNSW construction. Specifically, by leveraging the query process of C-HNSW, which will be introduced in Section 3.2.1, nodes are progressively inserted into the index starting from the top layer, connecting intra-layer links within the same layer and updating the inter-layer links. For lack of space, we give the details of the construction algorithm in Appendix A.2.
3.2. Online retrieval
In the online retrieval phase, after obtaining the query vector for a given question, ArchRAG generates the final answer by first conducting hierarchical search on the C-HNSW index, and then analyzing and filtering of retrieved information.
3.2.1. Hierarchical search.
We first introduce the query process of C-HNSW and then describe the hierarchical search. The query process is the core of C-HNSW, which can be used for constructing C-HNSW as illustrated in Algorithm 2.
input : $\mathcal{H}=({\mathcal{G}},{L_{inter}})$, $q$, $k$, $l$.
@@ 1 $s\leftarrow$ a random node in the highest layer $L$;
2 for $i\leftarrow L,\cdots,l+1$ do
3       $c\leftarrow$ SearchLayer(${G_{i}}=(V_{i},E_{i}),q,s,1,i$);
4       $s\leftarrow$ find the node in layer $i-1$ via the inter-layer links of $c$;
5
@@ 4       $s\leftarrow$ find the node in layer $i-1$ via the inter-layer links of $c$;
7 return $K$;
8Procedure SearchLayer(${G_{i}}=(V_{i},E_{i}),q,s,k,i$);
9 $V\leftarrow\{s\}$, $R\leftarrow\{s\}$, $Q\leftarrow$ initialize a queue containing $s$;
10 while $|R|>0$ do
@@ 11       $c\leftarrow$ nearest node in $Q$;
12       $f\leftarrow$ furthest node in $R$;
@@ 16             $V\leftarrow V\cup\{x\}$;
17             $f\leftarrow$ furthest node in $R$;
18             if $d(x,q)<d(f,q)$ or $|R|<k$  then
19                   $Q\leftarrow Q\cup\{x\}$, $R\leftarrow R\cup\{x\}$;
20                   if $|R|>k$ then remove $f$ from $R$;
21
return ${R}$;
Algorithm 2 C-HNSW query
Specifically, given a specified query layer $l$, query point $q$, and the number $k$ of nearest neighbors, the query algorithm can be implemented through the following iterative process:
  1. (1)
  2. (2)
    For each layer from layers $L$ to $l+1$, begin at the starting node and use a greedy traversal approach (i.e., the procedure SearchLayer) to find the nearest neighbor $c$ of $q$, and then traverse to the next layer using $c$’s inter-layer link as the starting node for the next layer (lines 2-4).
  3. (3)
    In the query layer $l$, use the greedy traversal approach to find the $k$ nearest neighbors of $q$ (line 5).
Specifically, the greedy traversal strategy compares the distance between the query point and the visited nodes during the search process. It achieves this by maintaining a candidate expansion queue $Q$ and a dynamic nearest neighbor set $R$, which contains $k$ elements:
@@ 19                   $Q\leftarrow Q\cup\{x\}$, $R\leftarrow R\cup\{x\}$;
  • Dynamic Nearest Neighbor Set $R$: Nodes added to $C$ are used to update $R$, ensuring that it maintains no more than $k$ elements, where $k$ is the specified number of query results.
In the greedy traversal, if a node $x$ expanded from $Q$ satisfies $d(n,q)>d(n,f)$, where $f$ is the furthest node from $R$ to $q$, then the traversal stops.
Based on the discussions above, we propose a hierarchical search approach, which retrieves the top $k$ relevant elements in each layer using the method outlined in Algorithm 2. Once the query at a given layer is complete, we use the inter-layer link of the nearest neighbor in that layer to traverse to the next layer, repeating the process until reaching the bottom layer. At the bottom layer, we extract the relationships between connected retrieved entities, forming the text information $R_{0}$ for this subgraph. We represent the retrieved information from each layer as $R_{i}$, where $i\in\{0,1,\cdots,L\}$, to be used in the adaptive filtering-based generation process.
3.2.2. Adaptive filtering-based generation.
While some optimized LLMs support longer text inputs, they may still encounter issues such as the “lost in the middle” dilemma (Liu et al., 2024b). Thus, direct utilization of retrieved information comprising multiple text segments for LLM-based answer generation risks compromising output accuracy.
@@
(3) $A_{i}$ $=LLM(P_{filter}||R_{i})$
(4) $Output$ $=LLM(P_{merge}||Sort(\{A_{0},A_{1},\cdots,A_{n}\}))$
@@
We also analyze the complexity of ArchRAG in Appendix A.3.
4. Experiments
Our experiments aim to answer following research questions (RQs):
  • RQ1: How does ArchRAG perform compared with existing baselines on specific and abstract QA tasks?
  • RQ2: How efficient is the ArchRAG approach?
  • RQ3: What is the quality of our LLM-based attributed communities, and how does it impact the performance of RAG?
  • RQ4: How robust is ArchRAG?
4.1. Setup
Table 2. Datasets used in our experiments. BLEU, MET, and ROU denote BLEU-1, METEOT, and ROUGE-L F1.
@@ 4.1. Setup
Metrics Accuracy, Recall Accuracy, Recall BLEU, MET, ROU
@@
We consider both specific and abstract QA tasks. For the specific QA task, we use the following datasets: Multihop-RAG (Tang and Yang, 2024), HotpotQA (Yang et al., 2018), and NarrativeQA (Kočiskỳ et al., 2018), all of which are extensively utilized within the QA and Graph-based RAG research communities (Hu et al., 2022; Gutiérrez et al., 2024; Yao et al., 2022; Xu et al., 2024; Sarthi et al., 2024; Asai et al., 2023). Multihop-RAG is designed to evaluate retrieval and reasoning across news articles in various categories. HotpotQA is a QA dataset featuring natural, multi-hop questions. NarrativeQA is a dataset that comprises QA pairs based on the full texts of books and movie transcripts. Besides, we follow the GraphRAG (Edge et al., 2024) method for the abstract QA task. We reuse the Multihop-RAG corpus and prompt LLM to generate questions that convey a high-level understanding of dataset contents. The statistics of these datasets are reported in Table 2.
Table 3. Performance comparison of different methods across various datasets for solving specific QA tasks. The best and second-best results are marked in bold and underlined, respectively.
@@
Graph-based RAG HippoRAG 38.9 19.1 51.3 56.8 2.2 5.0 2.8
LightRAG-Low 44.1 25.1 34.1 41.8 4.5 8.7 6.6
LightRAG-High 48.5 28.7 25.6 33.3 4.4 8.1 6.1
LightRAG-Hybrid 50.3 30.3 35.6 43.3 5.0 9.4 7.0
GraphRAG-Local 40.1 23.8 29.7 35.5 3.9 3.3 3.5
GraphRAG-Global 45.9 28.4 33.5 42.6 2.9 3.7 3.3
@@
Baselines.
Our experiments consider three model configurations:
  • Inference-only: Using an LLM for direct question answering without any retrieval data, i.e., Zero-Shot and CoT (Kojima et al., 2022).
  • Retrieval-only: Retrieval models extract relevant chunks from all documents and use them as prompts for large models. We select strong and widely used retrieval models: BM25 (Robertson and Walker, 1994) and vanilla RAG.
  • Graph-based RAG: These methods leverage graph data during retrieval. We select HippoRAG (Gutiérrez et al., 2024), GraphRAG (Edge et al., 2024), and LightRAG (Guo et al., 2024). Particularly, GraphRAG has two versions, i.e., GraphRAG-Global and GraphRAG-Local, which use global and local search methods respectively. Similarly, LightRAG integrates local search, global search, and hybrid search, denoted by LightRAG-Low, LightRAG-High, and LightRAG-Hybrid, respectively.
In GraphRAG-Global, all communities below the selected level are first retrieved, and then the LLM is used to filter out irrelevant communities. This process can be viewed as utilizing the LLM as a retriever to find relevant communities within the corpus. According to the selected level of communities (Edge et al., 2024), GraphRAG-Global can be further categorized into C1 and C2, representing high-level and intermediate-level communities, respectively, with C2 as the default.
Metrics.
For the specific QA tasks, we use Accuracy and Recall to evaluate performance on the first two datasets based on whether gold answers are included in the generations instead of strictly requiring exact matching, following (Schick et al., 2024; Mallen et al., 2022; Asai et al., 2023). We also use the official metrics of BLEU, METEOR, and ROUGE-l F1 in the NarrativeQA dataset. For the abstract QA task, we follow prior work (Edge et al., 2024) and adopt a head-to-head comparison approach using an LLM evaluator (GPT-4o). LLMs have demonstrated strong capabilities as evaluators of natural language generation, often achieving state-of-the-art or competitive results when compared to human judgments (Wang et al., 2023b; Zheng et al., 2023). Overall, we utilize four evaluation dimensions: Comprehensiveness, Diversity, Empowerment, and Overall. Please see Appendix B.1 for more details on metrics.
Implementation details.
We mainly use Llama 3.1-8B (Dubey et al., 2024) as the default LLM model and use nomic-embed-text (Nussbaum et al., 2024) as the embedding model to encode all texts. We use KNN for graph augmentation and the weighted Leiden algorithm for community detection. For each retrieval item $k$, we search the same number of items at each layer, with $k=5$ as the default. Our codes are available at: https://anonymous.4open.science/r/H-CAR-AG-1E0B. Additional details are provided in Appendix B.2.
4.2. Overall results
@@
To answer the question RQ1, we compare our method with baseline methods in solving both specific and abstract QA tasks.
$\bullet$ Results of specific QA tasks. Table 3 reports the performance of each method on three datasets. Clearly, ArchRAG demonstrates a substantial performance advantage over other baseline methods on these datasets. On the Multihop-RAG dataset, HippoRAG performs worse than retrieval-only methods while outperforming retrieval-based methods on the HotpotQA dataset. This is mainly because, on the HotpotQA dataset, passages are segmented by the expert annotators; that is, passages can provide more concise information, whereas on Multihop-RAG, passages are segmented based on chunk size, which may cause the LLM to lose context and produce incorrect answers. The experimental results suggest that not all communities are suitable for specific QA tasks, as the GraphRAG-Global performs poorly. Furthermore, GraphRAG does not consider node attributes during clustering, which causes the community’s summary to become dispersed, making it difficult for the LLM to extract relevant information. Thus, we gain an interesting insight: LLM may not be a good retriever, but is a good analyzer.
@@
$\bullet$ Results of abstract QA tasks. We compare ArchRAG against baselines across four dimensions on the Multihop-RAG dataset. Here, we only compare the LightRAG-Hybrid method, as it represents the best version of LightRAG methods (Guo et al., 2024). As shown in Figure LABEL:fig:winmap, our method achieves comparable performance with GraphRAG-Global on the diversity and empowerment dimensions while significantly surpassing it on the comprehensive dimension. Overall, ArchRAG demonstrates outstanding performance in solving abstract QA tasks.
$\bullet$ Efficiency of ArchRAG. To answer the question RQ2, we compare the time cost and token usage of ArchRAG with those of other baseline methods and the index consumption against that of graph-based RAG methods. As shown in Figure LABEL:fig:query-efficiency, ArchRAG demonstrates significant time and cost efficiency for online queries. For example, on the HotpotQA dataset, ArchRAG reduces token usage by a factor of 250 compared to GraphRAG-Global, utilizing 5.1M tokens and 1,394M tokens, respectively.
@@
Besides, Figure LABEL:fig:index-efficiency shows the index construction time and token usage. The cost of building index for ArchRAG is similar to that of GraphRAG, but due to the need for community summarization, both take higher time cost and token usage than HippoRAG.
LABEL:eff_index
4.3. Detailed Analysis
To answer the question RQ3, we first evaluate the community qualities under our proposed LLM-based attributed clustering framework and then conduct an ablation study.
$\bullet$ Community quality of different clustering methods. We evaluate 4 different clustering methods with 2 graph augmentation methods. The clustering methods we choose are weighted Leiden, Spectral Clustering (Von Luxburg, 2007), SCAN (Xu et al., 2007), and node2vec (Grover and Leskovec, 2016) with KMeans. The graph augmentation methods we choose are the KNN algorithm, which computes the similarity between each node, and CODICIL (Ruan et al., 2013), which selects and adds the top similar edges to generate better clustering results. We utilize the Calinski-Harabasz (CH) Index (Caliński and Harabasz, 1974) and Cosine Similarity (Charikar, 2002), abbreviated as CHI and Sim respectively, to evaluate the quality of generated attributed communities. The higher CH Index and Cosine Similarity indicate a higher quality of the community. The details of evaluation metrics can be found in Appendix B.1. As shown in Figures LABEL:fig:ch and LABEL:fig:sim, both KNN and CODICIL help generate high-quality communities well.
LABEL:quality_ch
LABEL:quality_sim
$\bullet$ Ablation study. To evaluate the effect of the attributed community in ArchRAG, we design some variants of ArchRAG and conduct the ablation experiments.
  • Spec.: Replaces the Weighted Leiden community detection algorithm with spectral clustering.
  • Spec. (No Aug): Removes the graph augmentation and uses spectral clustering.
  • Leiden: Replaces our clustering framework with the Leiden algorithm.
@@ 4.3. Detailed Analysis
The first two variants replace the algorithm in the LLM-based hierarchical clustering framework, while the last three variants modify key components of ArchRAG: attributes, hierarchy, and communities. As shown in Table 4, the performance of ArchRAG on specific QA tasks decreases when each feature is removed, with the removal of the community component resulting in the most significant drop. Additionally, the Direct variant demonstrates that the adaptive filtering-based generation process can effectively extract relevant information from retrieved elements.
Table 4. Comparing the performance of different variants of ArchRAG on the specific QA tasks.

### v3

@@ 12 ${G}(V,E)\leftarrow$ build a new graph using $C$ and $E^{\prime}$;
14until TT=True;
return HC{HC};
@@
C-HNSW index.
Given a query, to efficiently identify the most relevant information from each layer of the hierarchical tree $\Delta$, a naive method is to build a vector database for the ACs in each layer, which is costly in both time and space. To tackle this issue, we propose to build a single hierarchical index for all the communities. Recall that the ACs in $\Delta$ form a tree structure, and the number of nodes decreases as the layer level increases. Since this tree structure is similar to the HNSW (Hierarchical Navigable Small World) index which is the most well-known index for efficient ANN search (2018), we propose to map entities and ACs of $\Delta$ into high-dimensional nodes, and then build a unified Community-based HNSW (C-HNSW) index for them.
$\bullet$ The structure of C-HNSW. Conceptually, the C-HNSW index is a list of simple graphs with links between them, denoted by $\mathcal{H}=({\mathcal{G}},{L_{inter}})$ with $\mathcal{\mathcal{G}}=\{G_{0}=(V_{0},E_{0}),G_{1}=(V_{1},E_{1}),\cdots,G_{L}=(V_{L},E_{L})\}$, where $G_{i}$ is a simple graph and each node of the simple graph corresponds to an attributed community or entity. The number $L$ of layers in $\mathcal{H}$ is the same as that of $\Delta$. Specifically, for each attributed community or entity in the $i$-th layer of $\Delta$, we map it to a high-dimensional node in the $i$-th layer of $\mathcal{H}$ by using a language model (e.g., nomic-embed-text (2024)).
We next establish two types of links between these high-dimensional nodes, i.e., intra-layer and inter-layer links:
  • Intra-layer links: These links exist between nodes in the same layers. Specifically, for each node in each layer, we link it to at least $M$ nearest neighbors within the same layer, where $M$ is a predefined value, and the nearest neighbors are determined according to a given distance metric $d$. Thus, all the intra-layer links are edges in all the simple graphs: $L_{intra}=\bigcup_{i=0}^{L}E_{i}$.
  • Inter-layer links: These links cross two adjacent layers. Specifically, we link each node in each layer to its nearest neighbor in the next layer. As a result, all the inter-layer links can be represented as $L_{inter}=\bigcup_{i=1}^{L}\{(v,\psi(v))|v\in V_{i},\psi(v)\in V_{i-1}\}$, where $\psi(\cdot):V_{i}\rightarrow V_{i-1}$ is the injective function that identifies the nearest neighbor of each node in the lower layer.
For example, in Figure 2, the C-HNSW index has three layers (simple graphs), incorporating all the ACs. Within each layer, each node is connected to its two nearest neighbors via intra-layer links, denoted by undirected edges. The inter-layer links are represented by arrows, e.g., the green community at layer $L_{1}$ is connected to the green entity (its nearest neighbor at layer $L_{0}$).
@@
$\bullet$ The construction of C-HNSW. We propose a top-down approach to build a C-HNSW. Specifically, by leveraging the query process of C-HNSW, which will be introduced in online retrieval, nodes are progressively inserted into the index starting from the top layer, connecting intra-layer links within the same layer and updating the inter-layer links. For lack of space, we give the details of the construction algorithm in the appendix.
Online retrieval
In the online retrieval phase, after obtaining the query vector for a given question, ArchRAG generates the final answer by first conducting a hierarchical search on the C-HNSW index and then analyzing and filtering the retrieved information.
Hierarchical search.
We propose an efficient and fast retrieval algorithm, hierarchical search, to retrieve nodes from each layer of the C-HNSW structure. Intuitively, retrieving nodes from a given layer in C-HNSW requires starting from the top layer and searching downward through two types of links (i.e., intra-layer and inter-layer links) to locate the nearest neighbors at the given layer. In contrast, our hierarchical search algorithm accelerates this process by reusing intermediate results, the nearest neighbors found in higher layers, as the starting node for lower layers. This approach avoids redundant computations that would otherwise arise from repeatedly searching from the top layer, thereby enabling efficient multi-layer retrieval.
Algorithm 2 illustrates hierarchical search. Given the C-HNSW $\mathcal{H}$, query point $q$, and the number $k$ of nearest neighbors to retrieve at each layer, the hierarchical search algorithm can be implemented by the following iterative process:
  1. 1.
  2. 2.
    For each layer $i$ from the top layer $L$ down to layer 0, the algorithm begins at the starting node and performs a greedy traversal (i.e., the SearchLayer procedure) to find the set $R_{i}$ of the $k$ nearest neighbors of $q$. The set $R_{i}$ is then merged into the final result set $R$ (lines 4–5).
  3. 3.
    The closest neighbor $c$ of $q$ is then obtained from $R_{i}$, and the algorithm proceeds to the next layer by traversing the inter-layer link of $c$, using it as the starting node for the subsequent search (lines 6–7).
input : $\mathcal{H}=({\mathcal{G}},{L_{inter}})$, $q$, $k$.
@@ 1 $s\leftarrow$ a random node in the highest layer $L$;
2 $R\leftarrow\emptyset$;
3 for $i\leftarrow L,\cdots,0$ do
4 $R_{i}\leftarrow$ SearchLayer (${G_{l}}=(V_{l},E_{l}),q,s,k$);
5 $R\leftarrow R\cup R_{i}$;
6 $c\leftarrow$ get the nearest node from $R_{i}$;
7 $s\leftarrow$ find the node in layer i1i-1 via the inter-layer links of $c$;
8
9return $R$;
@@ 7 $s\leftarrow$ find the node in layer i1i-1 via the inter-layer links of $c$;
11 $V\leftarrow\{s\}$, $K\leftarrow\{s\}$, $Q\leftarrow$ initialize a queue containing $s$;
12 while $|K|>0$ do
@@ 13    $c\leftarrow$ nearest node in $Q$;
14    $f\leftarrow$ furthest node in $K$;
@@ 18       $V\leftarrow V\cup\{x\}$;
19       $f\leftarrow$ furthest node in $K$;
20       if $d(x,q)<d(f,q)$ or $|K|<k$  then
21          $Q\leftarrow Q\cup\{x\}$, $K\leftarrow K\cup\{x\}$;
22          if $|K|>k$ then remove $f$ from $K$;
23
24 return K{K};
Algorithm 2 Hierarchical search
Specifically, the greedy traversal strategy compares the distance between the query point and the visited nodes during the search process. It maintains a candidate expansion queue $Q$ and a dynamic nearest neighbor set $K$ containing $k$ elements, along with a stopping condition:
@@ 21          $Q\leftarrow Q\cup\{x\}$, $K\leftarrow K\cup\{x\}$;
  • Dynamic Nearest Neighbor Set $K$: Nodes added to $C$ are used to update $K$, ensuring that it maintains no more than $k$ elements, where $k$ is the number of query results.
  • Stopping Condition: The traversal terminates if a node $x$ expanded from $Q$ satisfies $d(n,q)>d(n,f)$, where $f$ is the furthest node in $K$ from the query point $q$.
After completing the hierarchical search and obtaining the ACs and entities from each layer, we further extract their associated textual information. In particular, at the bottom layer, we also extract the relationships between the retrieved entities, resulting in the textual subgraph representation denoted as $R_{0}$. We denote all the retrieved textual information from each layer as $R_{i}$, where $i\in{0,1,\dots,L}$, which will be used in the adaptive filtering-based generation process.
Adaptive filtering-based generation.
While some optimized LLMs support longer text inputs, they may still encounter issues such as the “lost in the middle” dilemma (Liu et al. 2024b). Thus, direct utilization of retrieved information comprising multiple text segments for LLM-based answer generation risks compromising output accuracy.
@@
$A_{i}$ $=LLM(P_{filter}||R_{i})$ (1)
$Output$ $=LLM(P_{merge}||Sort(\{A_{0},A_{1},\cdots,A_{n}\}))$ (2)
@@
Experiments
In this section, we conduct a comprehensive evaluation of our ArchRAG, focusing on both efficiency and performance.
Setup
Table 1: Datasets used in our experiments. Acc, Rec, Blue, Met, and Rou denote Accuracy, Recall, BLEU-1, METEOR, and ROUGE-L F1.
@@
Metrics Acc, Rec Acc, Rec Blue, Met, Rou
@@
We evaluate ArchRAG on both specific and abstract QA tasks. For specific QA, we use Multihop-RAG (2024), HotpotQA (2018), and NarrativeQA (2018), all of which are extensively utilized within the QA and Graph-based RAG research communities (2022; 2024; 2022; 2024; 2024; 2023). For abstract QA, we follow the GraphRAG (2024) method and reuse the Multihop-RAG corpus, prompting LLM to generate questions that convey a high-level understanding of dataset contents. The statistics of these datasets are reported in Table 1.
Baselines.
Our experiments consider three configurations:
  • Inference-only: Using an LLM to answer questions without retrieval, i.e., Zero-Shot and CoT (2022).
  • Retrieval-only: Retrieval models extract relevant chunks from all documents and use them as prompts for LLMs. We select strong and widely used retrieval models: BM25 (1994) and Vanilla RAG.
  • Graph-based RAG: These methods leverage graph data during retrieval. We select RAPTOR (2024), HippoRAG (2024), GraphRAG (2024), and LightRAG (2024). Particularly, GraphRAG has two versions, i.e., GGraphRAG and LGraphRAG, which use global and local search methods, respectively. Similarly, LightRAG integrates local search, global search, and hybrid search, denoted by LLightRAG, HLightRAG, and HyLightRAG, respectively.
In GGraphRAG, all communities below the selected level are first retrieved, and then the LLM is used to filter out irrelevant communities. This process can be viewed as utilizing the LLM as a retriever to find relevant communities within the corpus. According to the selected level of communities (2024), GGraphRAG can be further categorized into C1 and C2, representing high-level and intermediate-level communities, respectively, with C2 as the default.
Metrics & Implementation.
For the specific QA tasks, we use Accuracy and Recall to evaluate performance on the first two datasets based on whether gold answers are included in the generations instead of strictly requiring exact matching, following (2024; 2022; 2023). We also use the official metrics of BLEU, METEOR, and ROUGE-l F1 in the NarrativeQA dataset. For the abstract QA task, we follow prior work (2024) and adopt a head-to-head comparison approach using an LLM evaluator (GPT-4o). Overall, we utilize four evaluation dimensions: Comprehensiveness, Diversity, Empowerment, and Overall. For implementation, we mainly use Llama 3.1-8B (2024) as the default LLM and use nomic-embed-text (2024) as the text embedding model. We use KNN for graph augmentation and the weighted Leiden algorithm for community detection. For retrieval item $k$, we search the same number of items at each layer, with $k=5$ as the default. All methods are required to complete index construction and query execution within 3 days, respectively. Additional details are provided in the appendix, and our codes are provided in the supplementary material.
Table 2: Performance comparison of different methods across various datasets for solving specific QA tasks. The best and second-best results are marked in bold and underlined. OOT: Not finished within 3 days.
@@
Graph-based RAG RAPTOR 59.1 34.1 N/A N/A 5.5 12.5 9.1
HippoRAG 38.9 19.1 51.3 56.8 2.2 5.0 2.8
LLightRAG 44.1 25.1 34.1 41.8 4.5 8.7 6.6
HLightRAG 48.5 28.7 25.6 33.3 4.4 8.1 6.1
HyLightRAG 50.3 30.3 35.6 43.3 5.0 9.4 7.0
LGraphRAG 40.1 23.8 29.7 35.5 3.9 3.3 3.5
GGraphRAG 45.9 28.4 33.5 42.6 OOT OOT OOT
@@
Overall results
@@
We compare our method with baseline methods in solving both abstract and specific QA tasks.
@@
$\bullet$ Results of abstract QA tasks. We compare ArchRAG against baselines across four dimensions on the Multihop-RAG dataset. For the LightRAG, we only compare the HyLightRAG method, as it represents the best version (2024). As shown in Figure LABEL:fig:winmap, GGraphRAG outperforms other baseline methods, while our method achieves comparable performance on the diversity and empowerment dimensions and significantly surpasses it on the comprehensive dimension. Overall, by leveraging ACs, ArchRAG demonstrates superior performance in addressing abstract QA tasks.
$\bullet$ Results of specific QA tasks. Table 2 reports the performance of each method on three datasets. Note that GGraphRAG fails to complete querying on the NarrativeQA dataset within the 3-day time limit. RAPTOR is unable to build the index on datasets like HotpotQA, which contains a large number of text chunks. Its Gaussian Mixture Model (GMM) clustering algorithm requires prohibitive computational time and suffers from non-termination issues during clustering. Clearly, ArchRAG demonstrates a substantial performance advantage over other baseline methods on these datasets. The experimental results suggest that not all communities are suitable for specific QA tasks, as the GGraphRAG performs poorly. Furthermore, GraphRAG does not consider node attributes during clustering, which causes the community’s summary to become dispersed, making it difficult for the LLM to extract relevant information from a large number of communities. Thus, we gain an interesting insight: LLM may not be a good retriever, but is a good analyzer. We further analyze the reasons behind the underperformance of each graph-based RAG method and support our claims with empirical evidence in the appendix.
@@
$\bullet$ Efficiency of ArchRAG. We compare the time cost and token usage of ArchRAG with those of other baseline methods. As shown in Figure LABEL:fig:query-efficiency, ArchRAG demonstrates significant time and cost efficiency for online queries. For example, token usage on the HotpotQA dataset is cut by $\bf 250\times$ with ArchRAG compared to GraphRAG-Global, from 1,394M tokens down to 5.1M tokens.
To further evaluate ArchRAG, we test the efficiency of hierarchical search, indexing performance, and effectiveness on an additional dataset, RAG-QA Arena  (2024). Results show that ArchRAG achieves up to $5.4\times$ faster retrieval than basic HNSW, maintains efficient indexing, and achieves state-of-the-art performance on the RAG-QA Arena dataset. Additional details are provided in the appendix.
Detailed Analysis
To better understand the effectiveness of our proposed ArchRAG, we perform the following ablation study and experiment with a GGraphRAG variant.
$\bullet$ Ablation study. To evaluate the contributions of different components, we design several ArchRAG variants and conduct ablation experiments. These include two modifications to the LLM-based hierarchical clustering framework, three targeting core design elements in ArchRAG—attributes, hierarchy, and communities—and one direct prompting variant, as detailed below:
  • Spec.: Spectral clustering instead of weighted Leiden.
  • Spec. (No Aug): Spectral clustering without graph augmentation.
  • Leiden: Replaces our clustering framework with Leiden.
@@
As shown in Table 3, the performance of ArchRAG on specific QA tasks decreases when each feature is removed, with the removal of the community component resulting in the most significant drop. Additionally, the direct variant demonstrates that the adaptive filtering-based generation process can effectively extract relevant information from retrieved elements.
Table 3: Comparing the performance of different variants of ArchRAG on the specific QA tasks. Acc and Rec denote Accuracy and Recall, respectively.